# #1#1#1# 1 (1)

### v1

@@ Abstract
Long-context reasoning requires accurately identifying relevant information in extensive, noisy input contexts. Previous research shows that using test-time learning to encode context directly into model parameters can effectively enable reasoning over noisy information. However, meta-learning methods for enabling test-time learning are prohibitively memory-intensive, preventing their application to long context settings. In this work, we propose PERK (Parameter Efficient Reasoning over Knowledge), a scalable approach for learning to encode long input contexts using gradient updates to a lightweight model adapter at test time. Specifically, PERK employs two nested optimization loops in a meta-training phase. The inner loop rapidly encodes contexts into a low-rank adapter (LoRA) that serves as a parameter-efficient memory module for the base model. Concurrently, the outer loop learns to use the updated adapter to accurately recall and reason over relevant information from the encoded long context. Our evaluations on several long-context reasoning tasks show that PERK significantly outperforms the standard prompt-based long-context baseline, achieving average absolute performance gains of up to 90% for smaller models (GPT-2) and up to 27% for our largest evaluated model, Qwen-2.5-0.5B. In general, PERK is more robust to reasoning complexity, length extrapolation, and the locations of relevant information in contexts. Finally, we show that while PERK is memory-intensive during training, it scales more efficiently at inference time than prompt-based long-context inference.
@@ 1 Introduction
Large language models (LLMs) struggle with identifying and reasoning over relevant information in long, potentially noisy, input contexts (i.e., long-context reasoning [2, 34, 36, 38, 65]), including modern models with substantial context windows [32, 35, 23] such as GPT-4 (claimed context window of 128k tokens [23]). These challenges are not surprising, as longer contexts contain more irrelevant distractor information, and may also reflect more complex reasoning problems (i.e., multiple hops), both of which increase the challenge of reasoning tasks and degrade LLM performance [13]. Furthermore, positional biases in long-context models focus disproportionate attention on the beginning or end of contexts, neglecting intermediate information [39].
@@
In this work, we address these challenges of long-context reasoning by reframing it as test-time learning, enabling models to encode context at inference time using gradient-based parameter adaptation. Our approach processes and compresses lengthy sequences as a batch (or multiple batches) of shorter segments of the long context, rather than processing the full sequence token-by-token. The updated model can then use the newly stored parametric information (and its initial parametric knowledge) to answer relevant questions. During training, we operate two nested optimization loops: the inner loop, which is also used at test time, encodes the long context segment batches into model parameters, while the outer loop learns to retrieve and reason over relevant information from the updated parameters that now encode the long context.
@@
However, meta-learning algorithms for training test-time learning methods entail prohibitive memory overhead [17, 13, 25], particularly when updating the complete set of model parameters, because they require backpropagating through a complete multi-step optimization trajectory. To reduce the memory overhead, we introduce Parameter-Efficient Reasoning over Knowledge (PERK), a lighter-weight meta-learning algorithm that keeps training-time memory costs low by: (1) internalizing the context into a Low-Rank Adapter (LoRA [24]), and (2) backpropagating the outer loop over only a truncated trajectory from the final few inner-loop adaptation steps [54]. By lowering the memory cost, PERK scales more efficiently at training, supporting larger models and longer contexts.
@@
We benchmark PERK relative to models finetuned for classical in-context reasoning (FT-ICR) on multiple Needle-in-a-Haystack benchmarks, demonstrating PERK’s advantages. A small GPT-2 model (with 127M parameters and a native 1k context window) can be scaled by PERK to reason effectively over 8k token contexts, achieving a 60% absolute performance gain over fine-tuning GPT-2 for in-context reasoning. For larger models (Qwen-2.5-0.5B) with a native 32k context window, PERK surpasses FT-ICR by a 27% absolute improvement. As relevant facts in Needle-in-a-Haystack problems are often easily distinguishable from distractor text, we also propose a more challenging evaluation setting where relevant facts are surrounded by distributionally-similar distractor facts (i.e., Drops-in-the-Ocean). In this more complex setting, PERK models again outperform all baselines, including FT-ICR across the GPT-2, Qwen, and Mamba model families.
Further analyses reveal PERK’s strong test-time length extrapolation ability. Models trained with PERK on 1k-token contexts successfully extrapolate performance to unseen 32k-token contexts with only a 42% degradation, far better than FT-ICR’s 52% decline. In addition, unlike FT-ICR’s dramatic drops in performance (up to 90%) when relevant information shifts positions at test time, PERK robustly generalizes across contexts where relevant information is variably distributed along context positions. Finally, our inference complexity analysis shows that PERK’s memory footprint scales more efficiently than FT-ICR, as does its wall-clock runtime on extremely long contexts (> 32k tokens).
@@ 2 Background and Preliminaries
Causal Language Models   In this work, we focus on causal language models (CLM) $f_{\bm{\theta}}$ with a decoder-only transformer [51] architecture $f$ and parameters ${\bm{\theta}}$. The common objective in autoregressive-style training of a CLM is the token-level negative log-likelihood (NLL). For a collection of text sequences ${\mathcal{D}}=\{x\}$ with total length $N_{\mathcal{D}}$, this is defined as:
@@
For training, ${\bm{\theta}}$ is optimized iteratively via an optimization algorithm $\mathcal{A}lg:(\mathcal{L},{\mathcal{D}},{\bm{\theta}},\bm{h})\mapsto{\bm{\theta}}^{\prime}$ where $\mathcal{L}$ is the objective function, ${\mathcal{D}}$ is the training corpus, ${\bm{\theta}}$ are the initial parameters, ${\bm{\theta}}^{\prime}$ are the obtained parameter values, and $\bm{h}$ are the hyperparameters, such as optimizer choice and training length.
Reasoning with CLMs   We denote reasoning problems as tuples $r=({\mathcal{K}},{\bm{q}},y)$ of context ${\mathcal{K}}$, question ${\bm{q}}$, and target response $y$. With the pre-trained large CLMs’s ability to reason over facts in input contexts with reasonable success [14], prompt-based in-context reasoning (ICR) has become popular for automating reasoning tasks. In this setting, ${\mathcal{K}}$ and ${\bm{q}}$ are combined via a template into a single prompt ${\bm{x}}(r)$. The model’s parameters are optimized to predict the tokens of the correct answer:
@@
Reasoning as Test-time Learning   In this setting, models simulate reasoning by first encoding the given context ${\mathcal{K}}$ into a CLM’s parameters ${\bm{\theta}}$, and then using the updated parameters $\phi({\bm{\theta}},{\mathcal{K}})$ to generate a response to the given question ${\bm{q}}$ [13]. The context is encoded by adapting the model’s parameters using a CLM objective over the context: $\phi({\bm{\theta}},{\mathcal{K}})=\mathcal{A}lg(\mathcal{L}_{\text{NLL}},{\mathcal{K}},{\bm{\theta}},\bm{h})$. We then use $f_{\phi}$ to generate answers for any relevant questions ${\bm{q}}$.
Test-time learning methods for reasoning stem from a key observation: CLMs can store and retrieve diverse information in their parameters from pre-training [52, 5]. Yet these same models struggle with long-context problems [23, 32, 65, 35], despite the contexts containing far less information than what they’ve seen in pre-training. CLMs may reason better with knowledge in their parameters than in context, motivating methods that encode new context parametrically for reasoning.
Training a CLM for Test-time Learning   Given a distribution $R$ of possible reasoning problems, the optimal parameters ${\bm{\theta}}^{*}$ for a test-time learning CLM minimize the expected reasoning loss of the CLM’s adaptation to each reasoning problem $r\sim R$:
@@
Meta-learning algorithms, such as MAML [17], can be used for training effective test-time learning methods. In MAML, we obtain a meta-model $f_{{\bm{\theta}}}$ via bi-level optimization: the outer loop samples batches of reasoning problems from a training dataset of reasoning problems, the inner loop obtains the adaptations of ${\bm{\theta}}$ to each problem’s context ${\mathcal{K}}$, and the outer loop then optimizes the meta-parameters based on the performance of each $f_{\phi({\bm{\theta}},{\mathcal{K}})}$ on the corresponding questions. The outer loop optimization backpropagates through the entire inner loop, an operation that becomes increasingly expensive with more parameters and inner loop update steps. This approach—adapting a model with a CLM objective on context in the inner loop while meta-optimizing a reasoning objective in the outer loop—was first introduced by Reckoning [13], which demonstrated its potential.
@@ 3 PERK: Meta-Learning Parameter-Efficient Reasoning over Knowledge
While prior work has demonstrated the potential of test-time learning as a reasoning paradigm [13], these approaches are hindered by high memory overhead during training. The underlying meta-learning approach, MAML [17], requires backpropagating through the expected test-time learning phase (the inner loop). This gradient unrolling process involves computing higher-order derivatives, which are known to have substantial memory requirements [57], thereby limiting the number of model parameters that can be optimized and the effective sequence length that can be encoded.
To scale the training of larger and more capable test-time learning models, we propose PERK  (Parameter Efficient Reasoning over Knowledge), which uses a parameter-efficient test-time adaptation to reduce both the size and amount of gradient unrolling required during optimization. Specifically, in the inner and outer loops, we only optimize a parameter-efficient adapter, LoRA [24], to encode the contexts, and we apply truncated gradient unrolling [54] to reduce the memory cost of backpropagating through the inner loop. We visualize the PERK training process in Figure 1.
@@ 3.1 Parameter-Efficient Meta-Learning with Low-Rank Adaptation
Similar to Reckoning [13], PERK uses bi-level optimization to prepare a model for reasoning via test-time adaptation to any given context. At inference time, a model is adapted to the context via CLM optimization on that context, after which the adapted model is used to answer any relevant questions. However, unlike Reckoning, which directly finetunes the full model at training and test time, PERK only adapts a low-rank adapter [24]. During training, PERK encodes a context into the adapter through gradient-based updates, leaving the base model parameters unchanged. We denote by ${\bm{\theta}}_{\text{base}}$ and ${\bm{\theta}}_{\text{adapter}}$ the parameters of the base model and the LoRA adapter. ${\bm{\theta}}_{\text{base}}$ is never optimized, while ${\bm{\theta}}_{\text{adapter}}$ is meta-learned to a good initial state for its eventual test-time adaptations.
Test-Time Learning and Inner Loop Training    We set the test time learning objective (thus also the inner loop objective) to causal language modeling of the context: $\mathcal{L}_{\text{NLL}}({\mathcal{K}},({\bm{\theta}}_{\text{base}},{\bm{\theta}}_{\text{adapter}}))$, and set the adaptation algorithm $\mathcal{A}lg(\mathcal{L}_{\text{NLL}},{\mathcal{K}},({\bm{\theta}}_{\text{base}},{\bm{\theta}}_{\text{adapter}}),\bm{h})$ to only update ${\bm{\theta}}_{\text{adapter}}$. The adaptation algorithm computes the gradient $\nabla\mathcal{L}_{\text{NLL}}$ on a batch of sub-sequences from the full context sequence ${\mathcal{K}}$. This compression as a parallel batch (or multiple batches) allows us to process lengthy sequences beyond the model’s context window. The exact manner by which ${\mathcal{K}}$ is broken into batches is a hyperparameter of the algorithm, and for efficiency reasons, we set it to one batch during training.
Outer Loop Training    In the outer loop, we optimize the meta parameters ${\bm{\theta}}_{\text{adapter}}$ to learn from $R$, the corresponding distribution of reasoning problems. The optimal ${\bm{\theta}}_{\text{adapter}}$ minimize the expected reasoning loss of the adapter’s adaptation to each reasoning problem $r=({\mathcal{K}},{\bm{q}},y)\sim R$:
@@
where $\phi_{\text{adapter}}(({\bm{\theta}}_{\text{base}},{\bm{\theta}}_{\text{adapter}}),{\mathcal{K}})=\mathcal{A}lg(\mathcal{L}_{\text{NLL}},{\mathcal{K}},({\bm{\theta}}_{\text{base}},{\bm{\theta}}_{\text{adapter}}),\bm{h})$ is the parameter-efficient CLM adaptation of ${\bm{\theta}}_{\text{adapter}}$ to the context ${\mathcal{K}}$ obtained via the test-time adaptation. We may use any optimization algorithm to optimize ${\bm{\theta}}_{\text{adapter}}$, and in our experiments, we employ AdamW [40].
@@ 3.2 Scalable Meta-Learning with Truncated Gradient Unrolling
Gradient Unrolling (GU)    We denote an N-step inner-loop optimization’s intermediate states by:

### v2

# PERK: Long-Context Reasoning as Parameter-Efficient Test-Time Learning

@@ Abstract
Long-context reasoning requires accurately identifying relevant information in extensive, noisy input contexts. In this work, we propose PERK (Parameter Efficient Reasoning over Knowledge), a scalable approach for learning to encode long contexts using gradient updates at test time. Specifically, PERK employs two nested optimization loops in a meta-training phase. The inner loop rapidly encodes contexts into a low-rank adapter (LoRA) that serves as a parameter-efficient memory module for the base model. Concurrently, the outer loop learns to use the updated adapter to accurately recall and reason over relevant information from the encoded long context. Our evaluations on several long-context reasoning tasks show that PERK significantly outperforms the standard long-context finetuning, achieving average absolute performance gains of up to 20% for Qwen-2.5 (0.5B & 7B) on synthetic and real-world long-context reasoning. PERK also maintains its advantages across model scales and families. Compared to specialized long-context LLMs, PERK matches or surpasses their performance. Finally, our analyses show PERK is more robust to reasoning complexity, length extrapolation, and the positions of relevant information in contexts. https://perk-long-context.web.app
@@ 1 Introduction
Large language models (LLMs) struggle with identifying and reasoning over relevant information in long, potentially noisy, input contexts (i.e., long-context reasoning; anil2022exploring; lee2024longcontextlanguagemodelssubsume; li2024longcontextllmsstruggleicl; liu2023exposing; xu2024lifelongicl), including modern models with substantial context windows (kuratov2024babilong; levy2024tasktokensimpactinput; hsieh2024rulerwhatsrealcontext). These challenges arise because longer contexts contain more irrelevant distractor information, and may also be attached to more complex reasoning problems (i.e., multiple hops), both of which degrade LLM performance (chen2023reckoning). Furthermore, these challenges are exacerbated by positional biases in long-context models, which focus attention on the beginning or end of contexts, neglecting intermediate information (liu2024lost).
@@
In this work, we address these challenges of long-context reasoning by reframing it as test-time learning, enabling models to encode context at inference time using gradient-based parameter adaptation. Our approach, Parameter-Efficient Reasoning over Knowledge (PERK), processes and compresses lengthy sequences as a batch (or multiple batches) of shorter segments of the long context, rather than processing the full sequence token-by-token. The updated model can then use the newly stored parametric information (and its initial parametric knowledge) to answer relevant questions. The original context is then discarded in this answering step as it is now encoded parametrically by the model. To train PERK, we operate two nested optimization loops: the inner loop, which is also used at test time, encodes the long context segment batches into model parameters, while the outer loop learns to respond to queries using the updated parameters representing the context.
@@
However, bi-level optimization algorithms for training test-time learning methods entail prohibitive memory overhead (finn2017model; chen2023reckoning; hu2023onlineadaptation), particularly when updating all model parameters, because they require backpropagating through a complete multi-step optimization trajectory. To reduce memory overhead, PERK (1) internalizes the context into a Low-Rank Adapter (LoRA; hu2022lora) rather than the initial model parameters, and (2) backpropagates the outer loop over only a truncated trajectory from the final few inner-loop adaptation steps (truncated-pmlr-v89-shaban19a). Consequently, PERK scales efficiently at training, supporting larger models and longer contexts.
@@
We compare PERK to models finetuned on long contexts for classical in-context reasoning (FT-ICR) on multiple Needle-in-a-Haystack benchmarks, demonstrating PERK’s advantages. For Qwen-2.5-0.5B with a native 32K context window, PERK surpasses FT-ICR baselines by a 20% absolute improvement. PERK also outperforms specialized long-context LLMs (gao2025prolong; yang2025qwen251mtechnicalreport) with over 7B parameters that have undergone extensive long-context training, achieving the highest accuracy when generalizing to contexts up to 128k tokens. As relevant facts in Needle-in-a-Haystack problems are often easily distinguishable from distractor text, we also propose more challenging evaluation settings where relevant facts are surrounded by distributionally-similar distractor facts (e.g., Drops-in-the-Ocean, HotPotQA, TriviaQA). In these settings, PERK models again outperform the FT-ICR baselines, and even surpass commercial LLMs (e.g., GPT, Gemini) on benchmarks for long-context symbolic reasoning. PERK also works effectively across many model scales (0.1B, 0.5B, 8B) and families (GPT-2, Qwen-2.5, LLaMA-3.1/3.2).
Finally, our analysis also shows PERK’s advantages yield strong test-time length extrapolation. PERK trained on 8K-token contexts maintains decent performance on 64K and 128K windows, consistently surpassing FT-ICR length generalization, and even outperforming models with native 512K and 1M context windows. Finally, because PERK encodes sequences as a permutation-invariant batch of context window segments (rather than maintaining the full contiguous context), it generalizes across contexts where relevant information is variably distributed along context positions, unlike FT-ICR’s dramatic drops in performance (up to 90%) when relevant information shifts positions at test time.
@@ 2 Background and Preliminaries
Causal Language Models.Causal language models (CLM) $f_{\bm{\theta}}$ (Radford2018GPT) with parameters $\bm{\theta}$} are typically trained with an autoregressive learning objective such as the token-level negative log-likelihood (NLL). For a dataset ${\mathcal{D}}=\{x\}$ with $N_{\mathcal{D}}$ sequences, this loss is defined as:
@@
The initial parameters $\bm{\theta}$} are optimized iteratively via an optimization algorithm $\mathcal{A}lg:(\mathcal{L},{\mathcal{D}},{\bm{\theta}},\bm{h})\mapsto{\bm{\theta}}^{\prime}$, where $\bm{h}$ are hyperparameters, such as optimizer choice and training length.
Reasoning with CLMs.   We denote reasoning problems as tuples $r=({\mathcal{K}},{\bm{q}},y)$ of context $\mathcal{K}$}, question $\bm{q}$}, and target response $y$ (e.g., a question-answering dataset like HotpotQA; yang-etal-2018-hotpotqa). With the pretrained CLMs’s ability to reason over facts in input contexts with reasonable success (clark2020softreasoners), prompt-based in-context reasoning (ICR) has become popular for solving reasoning tasks. In this setting, $\mathcal{K}$} and $\bm{q}$} are combined into a single prompt ${\bm{x}}(r)$. The model’s parameters are optimized to predict the tokens of the correct answer:
@@
Reasoning as Test-time Learning.   In this setting, models simulate reasoning by first encoding the given context $\mathcal{K}$} into a CLM’s parameters $\bm{\theta}$}, and then using the updated parameters $\phi({\bm{\theta}},{\mathcal{K}})$ to generate a response to the given question $\bm{q}$} (chen2023reckoning). The context is encoded by adapting the model’s parameters using a CLM objective over the context: $\phi({\bm{\theta}},{\mathcal{K}})=\mathcal{A}lg(\mathcal{L}_{\text{NLL}},{\mathcal{K}},{\bm{\theta}},\bm{h})$. We then use $f_{\phi}$ to generate answers for any relevant questions $\bm{q}$}.
Test-time learning methods for reasoning stem from a key observation: CLMs can store and retrieve diverse information in their parameters from pre-training (roberts-etal-2020-much; bayazit-etal-2024-discovering). Yet these same models struggle with long-context problems (hsieh2024rulerwhatsrealcontext; kuratov2024babilong; xu2024lifelongicl; levy2024tasktokensimpactinput), despite the contexts containing far less information than what is seen during pre-training. CLMs may reason better with knowledge in their parameters than in context, motivating methods that encode new context parametrically for reasoning.
Training a CLM for Test-time Learning.   Given a distribution $R$ of possible reasoning problems, the optimal parameters $\bm{\theta}$$^{*}$ for a test-time learning CLM minimize the expected reasoning loss of the CLM’s adaptation to each reasoning problem $r\sim R$:
@@
Meta-learning algorithms, such as MAML (finn2017model), can be used for training effective test-time learning methods. In MAML, we obtain a meta-model $f_{{\bm{\theta}}}$ via bi-level optimization: the outer loop samples batches of reasoning problems from a training dataset of reasoning problems, the inner loop obtains the adaptations of $\bm{\theta}$} to each problem’s context $\mathcal{K}$}, and the outer loop then optimizes the meta-parameters based on the performance of each $f_{\phi({\bm{\theta}},{\mathcal{K}})}$ on the corresponding questions. The outer loop optimization backpropagates through the entire inner loop, an operation that becomes increasingly expensive with more parameters and inner loop update steps.
@@ 3 PERK: Meta-Learning Parameter-Efficient Reasoning over Knowledge
We propose PERK, a bi-level optimization algorithm for model reasoning via test-time adaptation for any given long context. At inference time, PERK adapts a base model to encode a context via CLM optimization on that context. Subsequently, the adapted model can be used to answer relevant questions about that context. To scale PERK’s adaptation to larger and more capable base LLMs for effective long-context reasoning, we use parameter-efficient methods to reduce both the size and amount of gradient unrolling required during optimization.
@@ 3.1 Parameter-Efficient Meta-Learning with Low-Rank Adaptation
During training, PERK optimizes a parameter-efficient adapter, LoRA (hu2022lora), instead of the complete model. Specifically, PERK encodes a context into the adapter through gradient-based updates, leaving the base model parameters unchanged. We denote by ${\bm{\theta}}_{\text{base}}$ and ${\bm{\theta}}_{\text{adapter}}$ the parameters of the base model and the LoRA adapter, respectively. ${\bm{\theta}}_{\text{base}}$ is never optimized, while ${\bm{\theta}}_{\text{adapter}}$ is meta-learned to a good initial state for test-time adaptations.
Test-Time Learning and Inner Loop Training.    We set the test time learning objective (thus also the inner loop objective) to causal language modeling of the context: $\mathcal{L}_{\text{NLL}}({\mathcal{K}},({\bm{\theta}}_{\text{base}},{\bm{\theta}}_{\text{adapter}}))$, and set the adaptation algorithm $\mathcal{A}lg(\mathcal{L}_{\text{NLL}},{\mathcal{K}},({\bm{\theta}}_{\text{base}},{\bm{\theta}}_{\text{adapter}}),\bm{h})$ to only update ${\bm{\theta}}_{\text{adapter}}$. The adaptation algorithm computes the gradient $\nabla\mathcal{L}_{\text{NLL}}$ on a batch of sub-sequences from the full context sequence $\mathcal{K}$}. This compression as a parallel batch (or multiple batches) allows us to process lengthy sequences beyond the model’s native context window. The exact manner by which $\mathcal{K}$} is broken into batches is a hyperparameter of the algorithm, and for efficiency reasons, we set it to one batch during training.
Outer Loop Training.    In the outer loop, the model learns to reason over the encoded information in the adapter (memory scratchpad) given a problem $\bm{q}$} without access to the long context $\mathcal{K}$} explicitly in the text space. We optimize the meta parameters ${\bm{\theta}}_{\text{adapter}}$ to learn from $R$, the corresponding distribution of reasoning problems. The optimal ${\bm{\theta}}_{\text{adapter}}$ minimize the expected reasoning loss of the adapter’s adaptation to each reasoning problem $r=({\mathcal{K}},{\bm{q}},y)\sim R$:
@@
where $\phi_{\text{adapter}}(({\bm{\theta}}_{\text{base}},{\bm{\theta}}_{\text{adapter}}),{\mathcal{K}})=\mathcal{A}lg(\mathcal{L}_{\text{NLL}},{\mathcal{K}},({\bm{\theta}}_{\text{base}},{\bm{\theta}}_{\text{adapter}}),\bm{h})$ is the parameter-efficient CLM adaptation of ${\bm{\theta}}_{\text{adapter}}$ to the context $\mathcal{K}$} obtained via the test-time adaptation. We may use any optimizer for ${\bm{\theta}}_{\text{adapter}}$, and in our experiments, we use AdamW (loshchilov2017decoupled). We describe further hyperparameters, training & inference details, and long-context data preprocessing in Appendix B.
@@ 3.2 Scalable Meta-Learning with Truncated Gradient Unrolling
Gradient Unrolling (GU).    We denote an N-step inner-loop optimization’s intermediate states by: